Journalist Report- March 22nd

This morning’s EVA (Extra Vehicular Activity) was a bit of a special one. Four of our crew members exited the vault, armed with a camera and a drone, and entered the 4×4 electrical rovers. They drove them south and stopped a bit further than the formation known as Zubrin’s Head. There, they exited the vehicles and walked west, towards the base of a single, conical, towering red mountain. At around 100m (300ft), it stood out as the tallest rock formation of the area, and they were ready to climb it. Despite sharply inclined slopes on either side, and protruding rocks near the top, the team was keen on making it to the top. With the drone, they circled the mountain, and determined an optimal path that went around the back side. On they went. This wasn’t the worst climb any of them had seen, but with the heavy space-suits on their backs, restricted vision, lack of water and limited supply of oxygen, it sure felt like one of the bad ones. Despite the difficulties, after half an hour of huff ‘n puff they finally reached the top, from which they could see the whole valley: snowy mountains far in the south, the Skyline Rim in the west, and the Martian plains stretching all the way to the horizon in the north, well past the MDRS and the area we have had the luck to explore these past two weeks.

This was their last EVA. On Saturday, they will embark on the shuttle which will take them away from this red land, away from the dry sand, the storms and the burning sun, and take them back home, back to Earth. Seeing, from above, the area they have spent these last two weeks in was a touching moment for them, and they felt the incoming goodbyes.

Back at base though, there were still things to do before departure. For starters, today was Michael’s birthday, and the whole team celebrated with birthday cake, board games and a custom-made paper tiara for Michael. The rest of the day was devoted to finishing up our experiments, and slowly beginning to clean up the base before the incoming crew gets here on Saturday. Indeed, there is only one day left for us here. And we’re trying to enjoy it as much as we can!

Greenhab Report – March 22nd

GreenHab report for Sol11

Crewmember’s name: Frédéric Peyrusson (Biologist)
Date 22/03/2018 (Mick’s birthday)

Environmental control:

Ambient with door open.
Shade cloth on.
Fan on, without cooling.

Average temperatures: 26°C
Low: 8°C
High: 32°C

Hours of supplemental light: 0 hours

Changes to crops: No change

Daily water usage for crops: 10 gal

Time(s) of watering for crops: 4

Moringa research observations: N/A

Changes to research plants: N/A

Mario finished his 3 hydroponic towers and transplanted lettuce and arugula. All other plants have been watered.

Support/supplies needed: N/A

Operations Report – 21st March

Crew 190 Operations Report 21th March 2018

SOL: 10

Name of person filing report: Bastien BAIX

Non-nominal systems: Two water pumps: tank-to-tank pump, interior Hab pump

Notes on non-nominal systems: tank-to-tank pump is taking significantly longer than normal (pump is used to transfer water from the exterior water tank to the interior tank). Interior pump, used to inject water into the Hab water system is making strange noises. Cause is unknown. Should we worry about it ?

Generator (hours run): Turned off at 10.40 am and turned on at … pm (not yet turned on).

Solar – SOC …% (Before generator is run at night)

Diesel – ~35%

Propane – 80%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~525 gallons

Water (third tank) – 0 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 529.4

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.7

Beginning charge: 100%

Ending charge: –

Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: 26.9

Beginning charge: 100%

Ending charge: 95%

Currently charging: YES

Opportunity rover used: YES

Hours: 25,7

Beginning charge: 100%

Ending charge: 90%

Currently charging: YES

Curiosity rover used: NO

Hours: 20.4

Beginning charge: 100%

Ending charge: –

Currently charging: NO

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 180 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Bastien and Martin working on their experiments

Summary of GreenHab operations: Mario and Fred working on their experiments

Summary of ScienceDome operations: Ariane, Michael, Maximilien and Fred working on their experiments

Summary of RAMM operations: Sophie working on her muon detector

Questions, concerns and requests to Mission Support: –

Journalist Report Sol 10 – 21st March

Slowly but surely, our stay at the MDRS is coming to an end. What seemed like a daunting two weeks at the beginning has passed right by us, and we are now only two days away from the moment our shuttle will take off from Mars and head back to Earth. Our experiments are slowly coming to an end, and we can all feel the end approaching.

Our commander, Maximilien, has been the cornerstone of our project since the beginning, a year ago, when we first began searching for funds and planning this very unique voyage. He has managed the team’s efforts with great precision, while at the same time giving more of himself than anyone else for the success of this project!

He is an industrial engineering graduate currently pursuing a PhD in organic synthesis (more specifically cyclopentane and 1,4-diene formation), and his experiment at the MDRS was centred on the characterization of physico-chemical properties of soil. His plan was to measure different characteristics of soils sampled at various locations on the Martian landscape, testing them for the factors which are most important for plant growth. These included pH, conductivity, B, total N, P, ions and macroscopic texture. Some of these factors are easy to measure with the basic instruments available locally, others with the spectrophotometer that we brought with us, and some needed more complex methods, requiring rare chemicals and solvents. These supplies were on the same shipment as Ariane’s culture media, which had an accident on its way to Mars and had to reverse directions and head back to Earth. This prevented Max from performing some of his planned experiments, but nonetheless, he has managed to isolate significant results with what was available locally. Still, this has freed up a bit of time for him, and so he has been able to share his extensive knowledge of chemistry to help others with their experiments – Mario to test the conductivity and pH of his hydroponic solutions, pH testing for Fred’s and Ariane’s bacteria, and calibration curves for Martin. As so, his time has been well used despite the unexpected outcome of his experiment!

Tomorrow, we will celebrate Michael’s birthday with an EVA in the morning, and some cake and board games or a movie in the evening! He sure would have liked some good beer on the side, but unfortunately there is a very strict policy on alcohol on Mars – it is completely forbidden! He will have to wait until we get back to Earth to enjoy a cold one..

GreenHab Report Sol 10 – 21st March

Crewmember’s name: Frédéric Peyrusson (Biologist)

Date 21/03/2018

Environmental control:

Ambient with door open.

Shade cloth on.

Fan on, without cooling.

Average temperatures: 30°C
Low: 8°C
High: 34°C

Hours of supplemental light: 0 hours

Changes to crops: No change

Daily water usage for crops: 6 gal

Time(s) of watering for crops: 3

Moringa research observations: N/A

Changes to research plants: N/A

Measurements and pictures of hydrogel supplemented plants and crops. Interestingly, hydrogels seem to confer protection to hydric stress. All other plants have been watered.

Support/supplies needed: N/A

Sol Summary – 21st March

Summary Title:
Embarrassingly intense Interview sessions

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 11, 08:00

Sol Activity Summary:
8:15 Medic inspection
8:25 Breakfast
9:40 EVA: Martin (drug degradation), Maximilien (soil sampling), Ariane (soil sampling), Bastien (cartography)
11:05 End of EVA – Crew meeting – Relaxation
13:00 Lunch
14:00 Scientific & academic work – Interviews (Ariane)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 11: See attached image

Notes: EVA for Mario (pictures), Sophie (must place her muon detector outside), Bastien (cartography), Michael (support). Scientific in the afternoon. Frederic has plenty of academic work to perform. Most of the projects should ideally be finished by sol 11, sol 12 (last simulation fay) being kept in backup.

Anomalies in work:

Weather:
Very cloudy in the morning, blue sky in late afternoon (allowing a one hour telescope session).

Crew Physical Status:
A bit tired.

EVA:
See EVA report.

Reports to be file:
Commander report
Journalist report
EVA#11 report
EVA#12 request for Sol11
Engineer report
Pictures

Support Requested:
None

Astronomy Report – 21st March

Astronomy Report
Name: Sophie Wuyckens Crew: 190
date: 3/21/2018

Sky Conditions: Cloudy all the day long but clear sky from 4:00 PM.

Wind Conditions: no wind

Observation Start Time: 16:30

Observation End Time: 17:30

Summary: Today Mick wanted to learn how to manipulate the telescope. Therefore we have benefited from a thinning to go the Musk observatory during one hour. It was a short observation but efficient. We took one beautiful shot.
Objects Viewed: dark spot on the Sun + small prominences

Problems Encountered: iCap had some difficulties to open at first. But we have finally managed to work with it.

Cheers,

Sophie Wuyckens

EVA#11 Report – 21st March

EVA Report:

EVA #11 Crew members: Martin ROUMAIN (EVA Leader), Bastien BAIX, Ariane SABLON, Maximilien RICHALD.

Location:

Along Cactus Road 1104: 521500E, 4252500N

Time:

Departure at 09:33 a.m.

End at 11:03 a.m. (instead of 12:00 p.m. as planned)

Duration:

1 hour 30 minutes

Narrative:

We left the few Martian animals (they really just look like black cows) who were blocking the main road. After a few minutes, they decided to let us go. At 0956 we arrived on Cactus Road (which was not really easy to find). We decided to go and explore a canyon by foot. Max and Ariane took samples for their respective experiments. At 1028, we were back to our vehicles. At 1040, we saw rain in the distance. As it was very cloudy, we decided to cancel the rest of the EVA and to immediately go back to the station. We arrived near the Station at 1055. We replaced the MDRS panel which fell because of the wind. We then headed back to the Hab.

Purposes:

Soil sampling for Maximilien

Sampling for Ariane’s bacteriology experiment

3D Mapping with Bastien’s drone

Check for Frédéric’s and Martin’s samples

Vehicles: 2 Rovers (Spirit and Opportunity)

9:33 a.m.: All crew members in the main airlock, ready for depressurization.

9:36 a.m.: Beginning of the EVA, took the rovers towards Cactus Road.

9:42 a.m.: Saw a few Martian animals on the main road who looked like black cows.

9:56 a.m.: Stopped on Cactus Road to explore a canyon by foot and harvest samples for Ariane’s and Maximilien’s experiments.

10:28 a.m.: Back on the road, took the rovers to go further down the road.

10:40 a.m.: Saw rain in the distance, we then decided to abort the EVA and go back to the the station.

10:55 a.m.: Back at the station, replaced the MDRS panel that fell because of the wind.

11:00 a.m.: All crew members in the main airlock.

11:03 a.m.: Back in the Hab, end of the EVA.